8-pack control system working group Meeting Minutes

18 March 2003

In attendance: R Chestnut, B Hall, K Kim, J Nelson, J Rock, S Smith

Notes by: J Nelson



-Need to check out readback with real signals, check archiving of variables (pwr, pmin, pmax, etc), and check the plots



-Steve found a small bug with the processing control variables DAC settings.  Ron fixed these.



DAC/ADC Issues

-Janice will double check the phase mapping.  Kukhee fixed the phases so they all map to +- 180.  You can enter a number outside that range, but it will be converted to a number inside that range.


All of these issues are for later:


-TBD: The trickle down of the SLED variables.  This was specified at a previous meeting (numbers on SLED panel need to trickle down to I/Q and A/P PVs).  For processing, we might want this feature functional sooner rather than later. 

-TBD: the trickle down of the variables on the power distribution panel as well.  There are 4 PVs there called Klystron Sled-in phases.  These need to trickle down to the phase DACs as well.  This phase is to be added to the channel phase angle (IQ calib panel) and to the timestep phase before going out to the DAC.



-Does 8-pack need a backup hardware trip circuit?  JN will ask NLCTA OPs

-This is on hold for now:  SWE will make a PV for a temporary DIO channel as a reset to Joe Frisch’s kludge RE Interlock.  They will send the name to Janice who will put a button on the panel with the pulse width and the klystron attenuation.  Its DIO channel assignment should be on Andrew’s previously sent list.  Also need a channel for status when it trips off.

-(no change) The following item was deferred for a few weeks:

LLRF + NLCTA Ops need to sort out what the LLRF enable and the mod off buttons do


SCP Configs

-LLRF folks need to check mini-dbedit files for TRBRs and report problems if any



-Software ready to go.  Need to check out with real signals: PVs, plots, and archiving

      some of the checkout was done, Bob still troubleshooting archiving

-This has been resolved:  there are now 8 installed in the crate (6 from before, 2 recently).  Kukhee reported the 6th of the 6 has the wrong trigger rate and the 2 recently installed need clock/trig setup, which Steve will provide.

-(no change) We deferred the following: longer term issue: what to do if reading out an SIS when a fault occurs?

-(no change) Right now the SISs are paused and read out with a user request only.  A read when a fault happens will also be added to this.



-Bob will update archive browser selection criteria per Janice’s email. 

-Janice will check current list of archived variables to see if any are missing.

(Notably PAVE, PMAX, PMIN, PWR and the processing control).

-The retrieval of archived waveforms is still problematic.  Bob is working on it.  He promises Matlab functions by the end of the week.


Archiving of SISs

-(no change) SWE plans to implement the following two features requested by LLRF (Joe’s email):

have waveforms archived on (1) every trip and (2) when an acquire is done by a user

-(no change) Later SWE will work with Joe on his other two requests: archive every few minutes and when a change occurs in a selected set of values in EPICS (a pseudo-trip if you will)



-nothing new 



(No change) All of the following will get going in the next few weeks:

-Each IQA channel has 6 or 7 calibration numbers that are module dependent.  Gain/offset I,Q, quadrature angle and overall phase.  These numbers are used to change the SIS waveforms from I/Q to A/P.  These numbers will go on the IQA module panel. 

-Steve will email the equations for this calculation

-Kukhee needs to change the SIS converter subrecord and new PVs need to be created.

-Once names are decided upon, Janice can make the panel (Judy may have started this already.)

-These 6 (or 7) numbers x 8 channels x 4 IQAs need to be archived at least once every 10 minutes.  The panel interface could just be one table for each IQA with 8 rows for the 8 channels and 6 columns with the calibration variables.


Processing algorithm

- Processing algorithm ready to be checked.  Steve, Kukhee and Janice will work on it on Friday.

-Steve needs to provide conversion from DAC volts to ns and dB this week.


Peak Power Meter (PPM)

-Plans for this week  (needed by Thursday 27th)

-The enet spigot has been allocated – it’s the first spigot after those labeled.  An enet box is in rack 113 level 45.  Tonee and Janice will run gpib cables to this box.

-(no change) Tonee will tweak the labview program to send the waveform to epics once we have names – probably late next week

-(no change) Tonee and Janice will set up the necessary hardware in ESB (Enet box, cables, etc)



-(no change) We’re currently waiting for Dick’s setup to settle down then we’ll do one more iteration of getting list, changing database, changing panels, programming the dcm, changing alarm handler and archiver stuff as needed



-Are we still confused with out vs sled?  Yes.  Janice will look.

-Is this resolved?  Steve will try to use Ron’s workaround for getting epics to printout on the color printer near his office.  Either with the vax queue name (SLC:Qname) or the NT queue name (something similar).

-(no change)Judy reported the crate profiles on the web are updated and look good.  Janice requested that channel assignment information also be available via the web, even it it’s just a link to the template file.  Judy and Ron will look into this.

-Has this been done?  Andrew will tell Kukhee when he takes out the DIO module to modify it for the deadman

-(no change) we deferred the following:

-public/private matlab issues to be resolved at a later date

-LLRF needs to decide what the reset button does.



Sort task list by person


Ron (and otherwise unassigned SWE tasks):

-(on hold for now)PV/database for DIO reset of Joe Frisch’s kludge RE Interlock

-trickle down of sled variables

-trickle down of phase variables

-database/PVs for 6 channels of calibration numbers for each IQA channel (also connect to Kukhee’s SIS converter subroutine)

-processing algorithm



-Processing algorithm



-checkout of waveform archiving, ADC readbacks, etc

-check list of variables being archived

-ask nlcta ops about hardware RE trip

-check out sled vs. out

-checkout phase wraparound

-panel for 6 calib numbers for each IQA channel

-Work with Tonee on PPM labview and hardware



-check into making module channel assignments available on the web as well as the crate profiles

-(maybe) update to-be-archived list generator to include changes mentioned above



-pass on values to convert DAC volts to ns and dB

-pass on equation for the 6 calibration numbers for each IQA channel

-(maybe) kludge interlock reset hardware



-matlab functions to fetch archived waveforms.

-fix paoutdata/cgi interface to archived waveforms.

-fix list of archived PVs

-fix archive browser chooser